OpenStudy (anonymous):

(a^4/-2b)^3 write your answer without parentheses

OpenStudy (anonymous):

@IQ250

OpenStudy (anonymous):

can anyone help me? please

OpenStudy (anonymous):

@TwiSTedSouls Can you help? ill give medal

OpenStudy (anonymous):

@TheSmartOne can you help me by guiding me? through it.

OpenStudy (anonymous):

@TTP anyway you can help me

TheSmartOne (thesmartone):

Is this your question? \[\Large\sf\left( \frac{ a^{-4} }{ -2b } \right)^3\]

OpenStudy (anonymous):

no the-4 is a 2 but everything else is correct

TheSmartOne (thesmartone):

If it is, we have to use this formula :) \(\sf\Large\left( \frac{a}{b}\right)^x = \frac{a^x}{b^x} \)

OpenStudy (anonymous):

\[(a ^{2}\div -2b)^{3}\] that's the problem so I would use that fomula ?

TheSmartOne (thesmartone):

So is your question... \(\Large\sf\left( \frac{ a^{2} }{ -2b } \right)^3\)

TheSmartOne (thesmartone):

Because you clearly have -4 in the question when you originally posted it .-.

Directrix (directrix):